mr chase purchased 300 adult and child tickets to the fair for a total of $4140.each adult ticket cost $15 and each child ticket cost $12.how many child tickets did he buy?
Answer by mananth(16946)

Adult x tickets
children y tickets
Equation for tickets

1 x + 1 y = 300 .............1

Equation for amount spent
15 x + 12.00 y = 4140 .............2
Eliminate y
multiply (1)by -12.00
Multiply (2) by 1
-12.00 x -12.00 y = -3600.00
15.00 + 12.00 y = 4140.00
Add the two equations
3 x = 540.00
/ 3
x = 180
plug value of x in (1)
180.00 x + 1.00 y = 300.00
180.00 + 1.00 y = 300.00
1.00 y = 120.00


Adult 180.00 tickets
children 120 tickets
m.ananth@hotmail.ca